Puff and Smoke agreed to share profits and losses in their partnership on a 7:3 basis, respectively, after a salary allowance of $25,000 is allocated to Puff. Earnings for the period total $115,000. What will be the total amount credited to Puff's Capital account when the Income Summary account is closed?

$90,000

$25,000

$88,000

$63,000

$52,000
Answer #1

Earnings after salary allowance

= 115,000 - 25,000

= 90,000

This is shared in the ratio of 7:3 among Puff and smoke

Amount credited to puff capital account

= 25,000 + (90,000*7/10)

= 25,000 + 63,000

= 88,000